### v2

# Merging Rate of Opinions via Optimal Transport on Random Measures

Abstract
Random measures provide flexible parameters for Bayesian nonparametric models. Given two different priors for a random measure, we develop a natural framework to investigate the rate at which the corresponding posteriors merge, as the sample size increases. We define a new distance between the laws of random measures that is built as a Wasserstein distance on the ground space of unbalanced measures, endowed with the bounded Lipschitz metric. We develop tight analytical bounds for its specification to completely random measures, including the special case of Poisson and gamma random measures. The bounds are interpreted in terms of an adapted extended Wasserstein distance between the Lévy measures and are used to investigate the merging between the posteriors of normalized gamma and generalized gamma priors. After a careful study on the identifiability of the law of the random measure, interesting asymptotic and finite-sample insights are derived without putting any assumption on the true data generating process.
Keywords. Bayesian nonparametrics; Completely random measures; Cox process; Impact of the prior; Lévy measure; Merging of opinions; Optimal transport; Poisson process; Wasserstein distance.
1 Introduction
In a Bayesian framework the distribution of the observations $(X_{i})_{i\geq 1}$ is usually modelled through a random parameter $\theta$. As the data arrives the prior distribution $\mathcal{L}(\theta)$ gets updated through conditional probability into $\mathcal{L}(\theta|X_{1},\dots,X_{n})$, known as the posterior distribution. A long-standing topic in Bayesian statistics is to understand and quantify the inferential impact of different priors on $\theta$. A natural way to proceed is to compute, for two different priors $\mathcal{L}(\theta^{1})$ and $\mathcal{L}(\theta^{2})$, a distance between their posteriors
$d(\mathcal{L}(\theta^{1}|X_{1},\dots,X_{n}),\mathcal{L}(\theta^{2}|X_{1},\dots,X_{n})).$
When the parameter space is finite-dimensional there are many well-studied proposals for a distance $d$. For example, assuming that $\theta$ is in a Euclidean space, Ley et al. (2017) have used the Wasserstein distance. This clearly does not cover Bayesian nonparametric (BNP) models, which use infinite-dimensional parameter spaces to increase their flexibility and avoid restrictive assumptions on the distribution of the data. In this work we focus on BNP models whose infinite-dimensional random parameter consists of a random measure $\tilde{\mu}$. The data on a Polish space $\mathbb{X}$ are modelled as conditionally independent and identically distributed (i.i.d.) given $\tilde{\mu}$, so that for every $n\geq 1$,
$X_{1},\dots,X_{n}|\tilde{\mu}\stackrel{{\scriptstyle\mbox{\scriptsize{\rm iid}}}}{{\sim}}t(\tilde{\mu}),$ (1)
for some transformation $t$. A popular choice for $t$ is the normalization $t(\tilde{\mu})=\tilde{\mu}/\tilde{\mu}(\mathbb{X})$, which defines a random probability measure whenever $0<\tilde{\mu}(\mathbb{X})<+\infty$ almost surely (Regazzini et al., 2003, James et al., 2006, James et al., 2009). As a classical example, Ferguson (1974) shows that the Dirichlet process is recovered by normalizing a gamma random measure. Other common choices include kernel mixtures for densities and hazard rates (Lo, 1984, Dykstra and Laud, 1981, James, 2005), exponential transformations for survival functions (Doksum, 1974) and cumulative transformations for cumulative hazards (Hjort, 1990), together with the extensive literature they generated.
Consider two different priors $\mathcal{L}(\tilde{\mu}^{1})$ and $\mathcal{L}(\tilde{\mu}^{2})$ for model (1). Our goal is to introduce a distance between posterior random measures,
$d(\mathcal{L}(\tilde{\mu}^{1}|X_{1},\dots,X_{n}),\mathcal{L}(\tilde{\mu}^{2}|X_{1},\dots,X_{n})),$ (2)
and use it to investigate three compelling questions that do not have a general answer in a non-dominated setting:
  • a.
    Given two different nonparametric priors, will the distance between the posteriors vanish as the number of observations goes to infinity, that is, will there be merging of opinions (Blackwell and Dubins, 1962)?
  • b.
    How fast does the distance vanish, that is, what is the merging rate?
  • c.
    Is the finite-sample distance strictly decreasing as a function of sample size, that is, does the merging start immediately or after a certain number of observations?
To this end, we devote the first part of this work to defining a natural distance between generic laws of random measures by relying on optimal transport (OT). OT studies the mass transportation problem between probability measures, and it is conveniently used to define distances between measures, the so-called Wasserstein distances, in terms of the minimal cost of transporting mass from one measure to another; see Villani (2009) and Santambrogio (2015) for exhaustive accounts. In principle, it defines a distance between probabilities on any Polish space. Our proposal is then to consider the Polish space of bounded measures endowed with the bounded Lipschitz (BL) distance, in a similar spirit to the Wasserstein over Wasserstein distance or Hierarchical Optimal Transport distance (Nguyen, 2016, Yurochkin et al., 2019, Dukler et al., 2019, Catalano and Lavenant, 2024). We study the topology of our Wasserstein over BL distance, showing that it metrizes the weak convergence over the weak convergence, together with the convergence of the total mass of the mean measure. The metrization of weak convergence is an extremely useful property in BNP (Nguyen, 2013, Nguyen, 2016, Ho et al., 2017) since one usually deals with (random) measures with atomic components and possibly disjoint supports.
The asymptotic analysis of the distance in (2) requires analytical upper and lower bounds. To this end, we focus our attention on a class of random measures that is usually used for models in (1). All the proposals we mentioned use completely random measures (CRMs) to define a prior on $\tilde{\mu}$. CRMs are a rich class of random measures defined by Kingman (1967), which includes Poisson, gamma, and generalized gamma random measures as specific examples. We refer to Lijoi and Prünster (2010) for an overview on CRMs as unifying thread of Bayesian nonparametric models. The law of a CRM without fixed atoms is characterized by a Lévy intensity $\nu$ on $(0,+\infty)\times\mathbb{X}$, which is a deterministic measure with potentially infinite mass, whose expression is usually very explicit and tractable. One of the core results of our work is to find a general upper bound of the Wasserstein over BL distance between CRMs in terms of an adapted discrepancy between the Lévy measures. We build this adapted discrepancy using the notion of extended Wasserstein distance $\mathcal{W}_{*}$ of (Figalli and Gigli, 2010, Guillen et al., 2019, Catalano et al., 2024), which remarkably allows for comparisons between measures with infinite mass, extending it to the multivariate space $(0,+\infty)\times\mathbb{X}$ in a similar spirit to the adapted (or nested) Wasserstein distance of Pflug and Pichler (2012), Bartl et al. (2024), and Backhoff et al. (2022). When restricted to random measures with the same mean measure, the adapted extended Wasserstein discrepancy provides a distance between Lévy intensities that is both analytically and computationally tractable, and dominates the Wasserstein over BL: this distance will play the main role in the second part of the work, which focuses on the use of the distance in Bayesian nonparametrics. We also extend its definition to random measures characterized by a random Lévy measure, such as Cox processes and Cox CRMs.
Before moving to the analysis of (2), one needs to investigate the identifiability of the random parameter $\tilde{\mu}$ in (1), that is, which random measures $\tilde{\mu}$ induce the same $t(\tilde{\mu})$. That is crucial to avoid outputting a strict positive distance even though the laws of the observation are the same, that is, to erroneously detect a lack of merging of opinion. Identifiability depends on the choice of transformation $t$ and the class of distributions for $\tilde{\mu}$: we focus on a specific setting, namely, normalization of CRMs. The task then amounts to understand which random measures induce the same normalization, divide the random measures in equivalence classes accordingly, and choose a representative for each equivalence class. This will lead to the definition of scaled CRMs $\tilde{\mu}_{\mathcal{S}}$, such that $\mathbb{E}(\tilde{\mu}_{\mathcal{S}}(\mathbb{X}))=1$, and their generalizations to Cox CRMs.
This sets the ground for a principled framework for studying the merging rate of opinions in terms of the distance between (the representatives of) the posteriors, as in (2). Relying on the almost-conjugacy of models (1) with respect to a CRM prior (James et al., 2009), we are able to shed new light on the use of the Dirichlet process and the normalized generalized gamma process, with unexpected results both from the asymptotic (Questions a. and b.) and finite-sample (Question c.) point of view.
We first focus on the Dirichlet process (Ferguson, 1973), which has been a cornerstone in the development of BNP, allowing one to make inference on the law of the observations through closed-form expressions. We use our setup to investigate to which extent different parameters for the Dirichlet process will lead to different learning outcomes, both in a finite sample and asymptotic scenario, without putting any assumption on the “true” data-generating process. In this context, we prove that the merging will always occur at a rate 1/$n$, where $n$ is the sample size, independently of the dimension of the space $\mathbb{X}$ and model’s parameters. Moreover, we show that if the two Dirichlet processes have the same concentration parameter the merging will start immediately with the first observation, whereas if they differ it is not always the case, even if the base probabilities $P_{0}$ coincide. From a statistical perspective this finding is quite remarkable: even if two Bayesians have the same prior guess $P_{0}$ of the distribution of the data, the difference between their opinions may increase if they see a finite amount of data, according to how much confidence they are willing to put on their prior guesses.
Next we move to generalized gamma CRMs (Brix, 1999), a versatile class of CRMs that recovers many well-known CRMs as special or limiting cases. This is often invoked as a generalization of the Dirichlet process that adds a discount parameter $\sigma$ to increase the flexibility of the model. Its normalization has been introduced by Lijoi et al. (2007) and used in many statistical analyses (e.g., De Blasi et al., 2009, Barrios et al., 2013, Griffin and Leisen, 2017, Caron and Fox, 2017, Lau and Cripps, 2022, Lee et al., 2023). We use our setup to investigate its merging with the Dirichlet process, again without putting any assumption on the data-generating process. We prove that the merging rate occurs at the rate $\max(n^{-1/(1+\sigma)},k/n)$, where $k$ is the number of distinct observations in the dataset. This result unravels the regimes where using a normalized generalized gamma CRM will actually guarantee a different learning outcome with respect to the Dirichlet process, with the following take-home message: either the number of observations $n$ is sufficiently small or the number of distinct values $k$ should be sufficiently large, and how large it must be depends on the choice of discount $\sigma$, namely, $k>n^{\frac{\sigma}{1+\sigma}}$.
Relation to previous work. This work is related to many different topics, including optimal transport on (completely) random measures, the merging of opinions, frequentist consistency and contraction rates, sensitivity and robustness analysis.
Optimal transport between random measures. The idea of endowing the space of measures, or probability measures, with a distance and then considering an optimal transport distance on laws of random measures with cost function being this distance between measures is not new. This has been done between laws of random probability measures when the distance is an optimal transport distance with recent applications in machine learning (Yurochkin et al., 2019, Dukler et al., 2019) and Bayesian statistics (Nguyen, 2016, Catalano and Lavenant, 2024). In a different direction, Huesmann (2016) extended this “OT over OT” problem to the case of random measures which are equivariant, which implies that the law of the random measures is invariant under translation. This corresponds to measures with infinite mass (because the space is unbounded), and the average transport cost is used as a measure of distance. Erbar et al. (2023) then studied the gradient flows in the geometry generated by this optimal transport problem. There have also been proposals of distances tailored for laws of point processes (see, e.g., Barbour and Brown, 1992, Schuhmacher and Xia, 2008, Decreusefond et al., 2016, and references therein), but in his work we look at random measures, not only point processes. In general terms, this forces us to consider also the distribution of the jumps of the random measures, and not only the one of their atoms.
Optimal transport on CRMs. The first proposals to define a distance between CRMs with tractable upper bounds can be found in Catalano et al. (2020, 2021), where the authors propose to use
$\sup_{A}\mathcal{W}(\mathcal{L}(\tilde{\mu}^{1}(A)),\mathcal{L}(\tilde{\mu}^{2}(A)))$
as a distance between CRMs, where $A$ is any Borel set and $\mathcal{W}$ is the Wasserstein distance on the Euclidean space. The key progress of Catalano et al. (2024) has been to define an OT distance directly at the level of the Lévy intensities, as follows. For a set $A$, $\mathcal{L}(\tilde{\mu}(A))$ is an infinitely divisible law with Lévy measure $\nu_{A}(\cdot)=\nu(\cdot,A)$. One can then use the extended Wasserstein distance $\mathcal{W}_{*}$ to define
$\sup_{A}\mathcal{W}_{*}(\nu_{A}^{1},\nu_{A}^{2}),$ (3)
which is a distance between CRMs (Catalano et al., 2024) and leads to closed-form expressions whenever i) the CRMs do not have fixed atoms; ii) the CRMs are almost surely (a.s.) finite and homogeneous, i.e., $\nu^{i}=\rho^{i}\otimes P^{i}$ is a product measure; iii) the CRMs share the same base measure, i.e., $P^{1}=P^{2}$. These conditions are usually met in the setting of Catalano et al. (2024) but are very restrictive in our framework since i) and ii) typically hold a priori but not a posteriori, and iii) entails that the atoms of the CRM have the same distribution, which is likely to be the first feature to change when choosing a different prior. Therefore we need to define the adapted extended Wasserstein discrepancy, a new OT distance on Lévy intensities with same mean measure that considers the joint distribution of the atoms and the jumps instead. Remarkably, it coincides with (3) whenever conditions i), ii), and iii) apply (see Remark 2 below).
Merging of opinions. The notion of merging of opinions was introduced and thoroughly studied in the acclaimed work of Blackwell and Dubins (1962), which holds for a very general setup beyond exchangeability. In principle their results cover model (1). However, in this setting, their assumptions can be considered restrictive for at least two reasons. First of all the merging is guaranteed for prior laws on $(X_{n})_{n\geq 1}$ that are absolutely continuous with one another. Whereas this holds in most parametric settings, it is not usually the case in nonparametric ones, where non-dominated models are the norm. As a classical example, two Dirichlet processes with different total base measures are mutually singular (Korwar and Hollander, 1973). It has also been observed by Diaconis and Freedman (1986), who avoid this condition by assessing the merging under a weaker topology. Moreover, both Blackwell and Dubins (1962) and Diaconis and Freedman (1986) assume that the data-generating mechanism coincides with one of the models under consideration, which may well not be the case in practice. Our aim is to overcome both restrictions and build a framework that does not require any assumption on the data. Indeed, we propose one of the few asymptotic analyses that does not put strong assumptions on the distribution of the data (e.g., to be a sampled from the model or i.i.d. from a true distribution): this is possible because we focus on the similarities between specific versions of the posteriors rather than on the adherence of the learning mechanism to the unknown truth.
Frequentist consistency and contraction rates. In the same spirit, our framework goes beyond frequentist consistency properties in that it does not assume the data to be i.i.d. from a true distribution. However, if one were willing to make this assumption, consistency of the model with respect to each prior would entail merging of opinions but would not provide meaningful information about the convergence rate. Indeed the contraction rate of each prior to the truth could be (and usually is) considerably slower than the merging rate. For example, in this work we show that two Dirichlet processes with different parameters always merge at the rate 1/$n$, independently of the dimension of the space. This is much faster than the parametric rate of $1/\sqrt{n}$, which is the best possible one for the contraction rate, and moreover it is well known that the latter deteriorates when the dimension increases. Thus, not only our framework has the advantage of not putting assumptions on the data-generating process, it also provides new prospectives when one is willing to put them.
Sensitivity analysis and robustness. To our knowledge, this work provides the first theoretical setup of global sensitivity analysis in Bayesian nonparametric non-dominated models, in the sense that it studies the performance of the model of interest with respect to any choice of prior. Previous studies on sensitivity for BNP models usually consist in varying the parameters of the priors or taking functional perturbations (Nieto-Barajas and Prünster, 2009) and providing empirical comparisons between the posteriors. In a dominated setting, more classical distances between (posterior) densities can be considered, as in Saha and Kurtek (2019). A recent work of Giordano et al. (2022) focuses on local sensitivity, which studies robustness with respect to small perturbation of the prior through differential approximations (Gustafson, 1996) and can be very useful when posterior distributions are not known in closed form. Since local sensitivity focuses on small perturbations, a good impact is always a small impact. On the contrary, our framework does not limit to small perturbations, and thus a big impact of the prior can also be regarded as a benign sign of flexibility of the model. Although our statistical analysis focuses on the latter aspect, our results may be read or adapted to treat robustness as well.
Structure of the paper. In Section 2 we define the Wasserstein over BL distance between the laws of random measures. We characterize its topology and provide a general lower bound, specializing it to Poisson random measures. In Section 3 we focus on CRMs and provide a universal upper bound in terms of the adapted extended Wasserstein discrepancy between their Lévy intensities. The bound is specialized to Poisson random measures and extended to Cox CRMs with a random Lévy intensity. In Section 4 we establish the class of (Cox) CRMs that are identifiable for model (1) with respect to normalization, leading to the notion of scaled (Cox) CRMs. The adapted extended Wasserstein discrepancy is a distance on this class: in Section 5 we study its finite sample and asymptotic behaviour to investigate the merging of opinions for BNP models. All proofs are deferred to the supplementary material. Table 1 summarizes the distances that we use in this work.
Table 1: Summary of the different distances we use in this work.
Symbol Name Space Introduced in
$\mathcal{W}_{d_{\mathbb{X}}}$ $1$-Wasserstein distance. $\mathcal{P}(\mathbb{X})$ Eq. (4)
$\mathcal{W}_{d_{\mathbb{X},t}}$ $1$-Wasserstein distance with metric $d_{\mathbb{X},\text{t}}$. $\mathcal{P}(\mathbb{X})$ Eqs. (4) and (6)
$\mathcal{W}_{*}$ Extended Wasserstein distance. $\mathcal{M}((0,+\infty))$ Def. 3
$\mathrm{BL}$ Bounded Lipschitz distance. $\mathcal{M}_{B}(\mathbb{X})$ Eq. (5)
$\mathcal{W}_{\mathrm{BL}}$ Wasserstein over bounded Lipschitz. $\mathcal{P}(\mathcal{M}_{B}(\mathbb{X}))$ Def. 1
$\mathrm{AD}$ Adapted extended Wasserstein discrepancy. $\mathcal{M}_{m}((0,+\infty)\times\mathbb{X})$ Def. 4 and Prop. 5
$\mathcal{W}_{\mathrm{AD}}$ Wasserstein distance over $\mathrm{AD}$. $\mathcal{P}(\mathcal{M}_{m}((0,+\infty)\times\mathbb{X}))$ Eq. (4) and Def. 4
2 Distance between measures and random measures
We start by discussing distances between probability measures and non-negative measures, with classical notions such as the Wasserstein distance and the bounded Lipschitz distance. Next, we define a distance between laws of random measures by lifting the distance between random measures into a distance between probabilities over random measures via optimal transport.
2.1 Distances between measures
We work on a Polish space $\mathbb{X}$ endowed with a distance $d_{\mathbb{X}}$ and its Borel $\sigma$-algebra $\mathcal{B}(\mathbb{X})$. We write $\mathcal{P}(\mathbb{X})$ for the set of probability measures, $\mathcal{M}(\mathbb{X})$ for the space of locally finite measures, and $\mathcal{M}_{B}(\mathbb{X})\subseteq\mathcal{M}(\mathbb{X})$ for the set of finite non-negative measures. The spaces $\mathcal{P}(\mathbb{X})$ and $\mathcal{M}_{B}(\mathbb{X})$ are endowed with the topology of weak convergence, making the maps $\mu\mapsto\int f\,\mathrm{d}\mu$ continuous for $f\in C_{b}(\mathbb{X})$ the space of continuous and bounded functions. It turns them into Polish spaces, and they are endowed with their Borel $\sigma$-algebra, which is the smallest one making the maps $\mu\mapsto\int f\,\mathrm{d}\mu$ measurable for $f\in C_{b}(\mathbb{X})$.
We recall that on a metric space $(\mathbb{Y},d_{\mathbb{Y}})$, given $p^{1},p^{2}$ two elements of $\mathcal{P}(\mathbb{Y})$, the optimal transport distance (of order $1$) between them is defined as
[…]
where the minimum is taken over $\Pi(p^{1},p^{2})$ the set of joint probability distributions having marginals $p^{1},p^{2}$, and the equality follows by Kantorovich duality. Here $1$-Lipschitz means $|f(x)-f(y)|\leq d_{\mathbb{Y}}(x,y)$ for all $x,y\in\mathbb{Y}$. In the sequel we will use this formula for $(\mathbb{Y},d_{\mathbb{Y}})=(\mathbb{X},d_{\mathbb{X}})$, but also when $\mathbb{Y}=\mathcal{M}_{B}(\mathbb{X})$ or $\mathbb{Y}=\mathcal{M}((0,+\infty)\times\mathbb{X})$ is a space of measures.
Restricting to the case $(\mathbb{Y},d_{\mathbb{Y}})=(\mathbb{X},d_{\mathbb{X}})$, note that with $\mathcal{W}_{d_{\mathbb{X}}}$ we cannot compare measures of different mass, that is, $\mathcal{W}_{d_{\mathbb{X}}}(p^{1},p^{2})$ becomes infinite if $p^{1}(\mathbb{X})\neq p^{2}(\mathbb{X})$ as can be seen by taking constant functions $f$. A natural generalization is the Bounded Lipschitz distance $\mathrm{BL}$, which replaces the supremum over $1$-Lipschitz functions in (4) by a supremum over $1$-bounded and $1$-Lipschitz functions. Specifically we define
[…]
where $1$-bounded means $|f(x)|\leq 1$ for all $x\in\mathbb{X}$. The distance $\mathrm{BL}$ metrizes the weak convergence on $\mathcal{M}_{B}(\mathbb{X})$ (Düll et al., 2021, Theorem 1.57). Regarding the comparison between $\mathcal{W}_{d_{\mathbb{X}}}$ and $\mathrm{BL}$ on $\mathcal{P}(\mathbb{X})$, note that we always have $\mathrm{BL}\leq\mathcal{W}_{d_{\mathbb{X}}}$. However the discrepancy can be large: $\mathrm{BL}$ is always bounded by $2$, while $\mathcal{W}_{d_{\mathbb{X}}}$ can be arbitrary very large if the first moments of the measures are large. To say something finer we introduce the “truncated” distance
$d_{\mathbb{X},\text{t}}=\min(d_{\mathbb{X}},2).$ (6)
This distance metrizes the same topology as $d_{\mathbb{X}}$ but is bounded. Moreover, every $1$-Lipschitz and $1$-bounded function is $1$-Lipschitz for $d_{\mathbb{X},\text{t}}$, while every $1$-Lipschitz function for $d_{\mathbb{X},\text{t}}$ is $1$-Lipschitz for $d_{\mathbb{X}}$ and $2$-bounded, so that $f/2$ is $1$-Lipschitz and $1$-bounded. Consequently, on $\mathcal{P}(\mathbb{X})$,
$\frac{1}{2}\mathcal{W}_{d_{\mathbb{X},\text{t}}}\leq\mathrm{BL}\leq\mathcal{W}_{d_{\mathbb{X},\text{t}}}\leq\mathcal{W}_{d_{\mathbb{X}}}.$ (7)
Thus $\mathcal{W}_{d_{\mathbb{X},\text{t}}}$ and $\mathrm{BL}$ are strongly equivalent metrics when restricted to $\mathcal{P}(\mathbb{X})$ but $\mathcal{W}_{d_{\mathbb{X},\text{t}}}$ easier to interpret as it corresponds to a classical Wasserstein distance, though with respect to a modified distance.
2.2 Distances between random measures
By a random measure $\tilde{\mu}$, we mean a measurable map from a probability space into $\mathcal{M}_{B}(\mathbb{X})$. As an example of random measure on $\mathbb{X}$, the reader could think of point processes such as Poisson random measures, or at completely random measures, which will be studied in the next section. The law of a random measure is an element of $\mathcal{P}(\mathcal{M}_{B}(\mathbb{X}))$. The latter space is endowed with the topology of weak convergence when $\mathcal{M}_{B}(\mathbb{X})$ is endowed with the topology of weak convergence, and we say that a sequence of random measures $(\tilde{\mu}_{n})_{n}$ converges in distribution to $\tilde{\mu}$ if the law of $\tilde{\mu}_{n}$ converges weakly to the one of $\tilde{\mu}$ as $n\to+\infty$. It is the case if and only if, for any $f\in C_{b}(\mathbb{X})$, the sequence of random variables $\int f\,\mathrm{d}\tilde{\mu}_{n}$ converges weakly to $\int f\,\mathrm{d}\tilde{\mu}$ as $n\to+\infty$ (Kallenberg, 2017, Theorem 4.19).
The formula (4) makes sense for any ground metric space $(\mathbb{Y},d_{\mathbb{Y}})$, and we will use it with $(\mathbb{Y},d_{\mathbb{Y}})=(\mathcal{M}_{B}(\mathbb{X}),\mathrm{BL})$. Indeed, interpreting laws of random measures as probability distributions over $\mathcal{M}_{B}(\mathbb{X})$, it is possible to lift the distance $\mathrm{BL}$ into a distance over the space $\mathcal{P}(\mathcal{M}_{B}(\mathbb{X}))$. We call the resulting distance the “Wasserstein over Bounded Lipschitz” distance, denoted by $\mathcal{W}_{\mathrm{BL}}$. If we were to look at laws of random probability measures, that is, elements of $\mathcal{P}(\mathcal{P}(\mathbb{X}))$, then the analogue of the Wasserstein over Bounded Lipschitz distance would be the “Wasserstein over Wasserstein” distance, which has been proposed and studied in different contexts (Nguyen, 2016, Yurochkin et al., 2019, Dukler et al., 2019, Catalano and Lavenant, 2024).
Definition 1.
If $\mathbb{Q}^{1}$, $\mathbb{Q}^{2}\in\mathcal{P}(\mathcal{M}_{B}(\mathbb{X}))$ are two laws of random measures, we define the distance $\mathcal{W}_{\mathrm{BL}}$ between them as
$\mathcal{W}_{\mathrm{BL}}(\mathbb{Q}^{1},\mathbb{Q}^{2})=\min_{\pi\in\Pi(\mathbb{Q}^{1},\mathbb{Q}^{2})}\mathbb{E}_{(\tilde{\mu}^{1},\tilde{\mu}^{2})\sim\pi}(\mathrm{BL}(\tilde{\mu}^{1},\tilde{\mu}^{2})).$
A random measure $\tilde{\mu}$ has finite mean if $\mathbb{E}(\tilde{\mu}(\mathbb{X}))<+\infty$. In this case its mean measure $\mathbb{E}(\tilde{\mu})$, defined by $\mathbb{E}(\tilde{\mu})(A)=\mathbb{E}(\tilde{\mu}(A))$, is an element of $\mathcal{M}_{B}(\mathbb{X})$.
Proposition 1.
The quantity $\mathcal{W}_{\mathrm{BL}}$ defines a complete distance over the laws of random measures with finite mean. If $(\tilde{\mu}_{n})_{n}$, and $\tilde{\mu}$ are random measures with finite mean, then $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}_{n}),\mathcal{L}(\tilde{\mu}))$ converges to 00 as $n\to+\infty$ if and only if $(\tilde{\mu}_{n})_{n}$ converges in distribution to $\tilde{\mu}$ and $\mathbb{E}(\tilde{\mu}_{n}(\mathbb{X}))$ converges to $\mathbb{E}(\tilde{\mu}(\mathbb{X}))$.
We also provide a universal lower bound of this distance. Meaningful upper bounds are delayed until the next section, when we assume more structure on the random measures.
Proposition 2.
For any random measures $\tilde{\mu}^{1},\tilde{\mu}^{2}$ with finite mean, there holds
$\mathrm{BL}(\mathbb{E}(\tilde{\mu}^{1}),\mathbb{E}(\tilde{\mu}^{2}))\leq\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}),\mathcal{L}(\tilde{\mu}^{2})).$
Example 1.
Let $\tilde{\mathcal{N}}^{1}$ and $\tilde{\mathcal{N}}^{2}$ be two Poisson random measures of intensities $\alpha_{i}P^{i}_{0}$, for $i=1,2$, where $\alpha_{i}>0$ and $P^{i}_{0}$ is a probability on $\mathbb{X}$. The lower bound of Proposition 2 yields
$\mathrm{BL}(\alpha_{1}P^{1}_{0},\alpha_{2}P^{2}_{0})\leq\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mathcal{N}}^{1}),\mathcal{L}(\tilde{\mathcal{N}}^{2})).$
In particular, taking $f=1$, we have $|\alpha_{1}-\alpha_{2}|\leq\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mathcal{N}}^{1}),\mathcal{L}(\tilde{\mathcal{N}}^{2}))$. On the other hand, if $\alpha_{1}=\alpha_{2}=\alpha$, then by (7) we have $\frac{\alpha}{2}\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P^{1}_{0},P^{2}_{0})\leq\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mathcal{N}}^{1}),\mathcal{L}(\tilde{\mathcal{N}}^{2}))$.
3 Completely random measures and bounds via optimal transport on Lévy intensities
In our applications to Bayesian statistics, the random measures are not arbitrary but rather possess some additional structure that will lead to tractable computations. In this section we focus on completely random measures (CRMs), a key ingredient to building many Bayesian nonparametric priors, and refer to Lijoi and Prünster (2010) for an overview. We derive a tractable upper bound on the $\mathcal{W}_{\mathrm{BL}}$ distance between CRMs and Cox CRMs. These upper bounds will be characterized in terms of an optimal transport distance on Lévy intensities, which is interesting per se and will be further investigated in the next sections.
In the original definition of Kingman (1967), completely random measures are random measures such that, for every disjoint $A_{1},\dots,A_{n}\in\mathcal{B}(\mathbb{X})$, the random variables $\tilde{\mu}(A_{1}),\dots,\tilde{\mu}(A_{n})$ are mutually independent. This definition entails that any CRM can be decomposed in the sum of a deterministic measure, a discrete random measure with fixed atoms and a discrete random measure without fixed atoms, whose law is characterized by a Lévy intensity. Since it is the third component that plays the main role in Bayesian nonparametrics, we will give a slightly more restrictive definition of completely random measure as follows.
Recall that $\mathcal{M}(\mathbb{Y})$ is the set of locally finite non-negative measures, possibly with infinite mass. Mostly we consider $\mathbb{Y}=(0,+\infty)$ or $\mathbb{Y}=(0,+\infty)\times\mathbb{X}$. Here and after, =$\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}$ denotes equality in distribution.
Definition 2.
A random measure $\tilde{\mu}$ is a completely random measure (CRM) if there exists $\nu\in\mathcal{M}((0,+\infty)\times\mathbb{X})$ with $\iint\min(1,s)\,\mathrm{d}\nu(s,x)<+\infty$ such that
$\tilde{\mu}(A)\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\iint_{(0,+\infty)\times A}s\,\mathrm{d}\tilde{\mathcal{N}}(s,x),$ (8)
where $\tilde{\mathcal{N}}$ is a Poisson random measure on $(0,+\infty)\times\mathbb{X}$ of intensity $\nu$. We write $\tilde{\mu}\sim\textup{CRM}(\nu)$ and refer to $\nu$ as the Lévy intensity of $\tilde{\mu}$.
Thanks to Campbell’s theorem, a CRM $\tilde{\mu}$ has a finite mean $\mathbb{E}(\tilde{\mu}(\mathbb{X}))<+\infty$ if and only if $M_{1}(\nu)=\iint s\,\mathrm{d}\nu(s,x)<+\infty$. Note that we do a slight abuse of notation, as $\nu$ is a measure on the product space $(0,+\infty)\times\mathbb{X}$, and by $M_{1}(\nu)$ we denote the first moment only with respect to the first coordinate.
Proposition 3.
Let $\tilde{\mu}\sim\textup{CRM}(\nu)$ with finite mean. Then $\nu$ can be uniquely decomposed as $\mathrm{d}\nu(s,x)=\mathrm{d}\rho_{x}(s)\mathrm{d}P_{0}(x)$, where $P_{0}$ is a probability measure and $\rho$ is a transition kernel on $\mathbb{X}\times\mathcal{B}(\mathbb{R}_{+})$ that satisfy, for $P_{0}$-almost every (a.e.) $x\in\mathbb{X}$,
$\int_{0}^{+\infty}s\,\mathrm{d}\rho_{x}(s)=\mathbb{E}(\tilde{\mu}(\mathbb{X}));\qquad P_{0}(\cdot)=\frac{\mathbb{E}(\tilde{\mu}(\cdot))}{\mathbb{E}(\tilde{\mu}(\mathbb{X}))}.$
We refer to $\mathrm{d}\nu(s,x)=\mathrm{d}\rho_{x}(s)\mathrm{d}P_{0}(x)$ as the canonical decomposition of $\nu$.
An important case is when $\rho_{x}$ does not depend on $x$. Typical densities used are $\rho(s)=\alpha s^{-1}\exp(-bs)$ or $\rho(s)=\alpha s^{-(1+\sigma)}\Gamma(1-\sigma)^{-1}\exp(-bs)$ corresponding to respectively a Gamma CRM and a generalized Gamma CRM, where $\alpha,b$ and $\sigma>0$ are parameters.
Remark 1.
Intuitively, the measure $P_{0}$ describes the distribution of atoms, while $\rho_{x}$ determines the conditional distribution of the jump size at location $x$. This can be understood by approximating (8) with a compound distribution, so that for every $\epsilon>0$ and set $A$,
$\iint_{(\varepsilon,+\infty)\times A}s\,\mathrm{d}\tilde{\mathcal{N}}(s,x)\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\sum_{i=0}^{N}s_{i}\,\delta_{x_{i}},$ (9)
where by letting $r=\nu((\epsilon,+\infty)\times A)$, $N$ is a Poisson random variable with mean $r$, and $(s_{i},x_{i})\sim P$ independently, with $P=\nu|_{(\epsilon,+\infty)\times A}/r$ being the normalized restriction of $\nu$ to $(\epsilon,+\infty)\times A$.
We move on to the definition of a distance at the level of Lévy intensities, which is actually related to the distance $\mathcal{W}_{\mathrm{BL}}$ defined in the previous section. The underlying idea is simple: if we can find a coupling between the Lévy intensities $\nu^{1},\nu^{2}$ and if $\tilde{\mu}^{i}\sim\textup{CRM}(\nu^{i})$ for $i=1,2$, then we have a coupling between $\tilde{\mu}^{1}$ and $\tilde{\mu}^{2}$, which means we should be able to find an upper bound of $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}),\mathcal{L}(\tilde{\mu}^{2}))$. The challenge is to understand how $\mathbb{E}(\mathrm{BL}(\tilde{\mu}^{1},\tilde{\mu}^{2}))$ relates to a quantity which depends only on the Lévy intensities.
This leads to the definition of a discrepancy between Lévy intensities $\mathrm{d}\nu(s,x)=\mathrm{d}\rho_{x}(s)\mathrm{d}P_{0}(x)$, which is ultimately justified by Theorem 4 below. We observe that the two components $\rho_{x}$ and $P_{0}$ play very different roles, as highlighted in Remark 1. Indeed, $P_{0}$ is a probability measure on $\mathbb{X}$ describing the law of the atoms, whereas $\rho_{x}$ is $P_{0}$-a.s. a measure with infinite mass on $(0,+\infty)$ that determines the conditional law of their corresponding jumps. We will thus preserve their different nature, in the spirit of the nested Wasserstein distance for probability measures, which was originally introduced by Pflug and Pichler (2012) and is often termed adapted Wasserstein distance in financial mathematics (Backhoff et al., 2022, Bartl et al., 2024). Specifically we impose that, if an atom $x$ is coupled to an atom $y$ on the space $\mathbb{X}$, then the whole measure $\rho_{x}$ is also coupled to $\rho_{y}$.
A key tool is a definition of an “extended” Wasserstein distance to compare measures of infinite mass $\rho_{x}$ and $\rho_{y}$ A (quadratic) Wasserstein distance between measures of different and potentially infinite mass has been proposed by Figalli and Gigli (2010) through the notion of extended coupling. We refer to Catalano et al. (2024) and Guillen et al. (2019) for a concise overview. In the present work, we will only consider measures on the half-real line $\mathbb{Y}=(0,+\infty)$, and look at Wasserstein distances of order $1$. We refer to Catalano et al. (2024, Proposition 4) for a link between the definition below and the previous works of Figalli and Gigli (2010) and Guillen et al. (2019).
Definition 3.
The extended Wasserstein distance (of order $1$) between two positive Borel measures $\rho^{1},\rho^{2}$ on $(0,+\infty)$ with finite first moments about 00 is defined as
$\mathcal{W}_{*}(\rho^{1},\rho^{2})=\int_{0}^{+\infty}|U_{1}(s)-U_{2}(s)|\mathrm{d}s,$ (10)
where, for $i=1,2$ and $s>0$, $U_{i}(s)=\rho^{i}(s,+\infty)$ is the tail integral of $\rho^{i}$.
Following the aforementioned literature, and using the extended Wasserstein distance, we now present what we call the “adapated extended Wasserstein discrepancy”.
Definition 4.
Let $\nu^{1},\nu^{2}$ be Lévy intensities with $M_{1}(\nu^{i})=\iint s\,\mathrm{d}\nu^{i}(s,x)<+\infty$ and canonical decomposition $\mathrm{d}\nu^{i}(s,x)=\mathrm{d}\rho^{i}_{x}(s)\mathrm{d}P^{i}_{0}(x)$ for $i=1,2$. We define the adapted extended Wasserstein discrepancy between $\nu^{1}$ and $\nu^{2}$ as
$\mathrm{AD}(\nu^{1},\nu^{2})=\inf_{\pi\in\Pi(P_{0}^{1},P_{0}^{2})}\;\mathbb{E}_{(X,Y)\sim\pi}\left(\frac{M_{1}(\nu^{1})+M_{1}(\nu^{2})}{2}d_{\mathbb{X},\text{t}}(X,Y)+\mathcal{W}_{*}(\rho_{X}^{1},\rho_{Y}^{2})\right).$
The presence of the first moments $M_{1}(\nu^{1})$ and $M_{1}(\nu^{2})$ may seem counterintuitive, and prevents $\mathrm{AD}$ from being a distance. It is however necessary in order for the next theorem to hold. Note also that we use the truncated distance $d_{\mathbb{X},\text{t}}$ instead of the original distance $d_{\mathbb{X}}$. This is again because of the next theorem as it gives a sharper upper bound; see also (7) and the discussion related to it to understand why $d_{\mathbb{X},\text{t}}$ is a natural distance to consider. Thanks to the use of this truncated distance, note that there is no finiteness of any moment of $P_{0}^{i}$, $i=1,2$ in the definition of $\mathrm{AD}$. Actually, from Fubini $\int_{0}^{+\infty}U_{\rho}(s)\,\mathrm{d}s=\int_{0}^{+\infty}s\rho(s)\mathrm{d}s$ and thus $\mathcal{W}_{*}(\rho^{1}_{X},\rho^{2}_{X})\leq M_{1}(\nu^{1})+M_{2}(\nu^{2})$. It yields, together with the boundedness of $d_{\mathbb{X},\text{t}}$, that
$\mathrm{AD}(\nu^{1},\nu^{2})\leq 2(M_{1}(\nu^{1})+M_{1}(\nu^{2})).$ (11)
We move to a key theorem justifying the introduction of this adapted extended Wasserstein discrepancy.
Theorem 4.
If $\tilde{\mu}^{i}\sim\textup{CRM}(\nu^{i})$ for $i=1,2$ have finite mean, then
$\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}),\mathcal{L}(\tilde{\mu}^{2}))\leq\mathrm{AD}(\nu^{1},\nu^{2}).$
Remark 2.
The adapted extended Wasserstein discrepancy $\mathrm{AD}$ is particularly convenient when at least one of the two Lévy intensities is homogeneous, i.e., $\nu^{1}=\rho^{1}\otimes P_{0}^{1}$. In such case, $\mathcal{W}_{*}(\rho_{X}^{1},\rho_{Y}^{2})^{p}=\mathcal{W}_{*}(\rho^{1},\rho_{Y}^{2})^{p}$ does not depend on the coupling and thus
$\mathrm{AD}(\nu^{1},\nu^{2})=\frac{M_{1}(\nu^{1})+M_{1}(\nu^{2})}{2}\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0}^{1},P_{0}^{2})+\mathbb{E}_{Y\sim P_{0}^{2}}(\mathcal{W}_{*}(\rho^{1},\rho_{Y}^{2})).$
When both Lévy intensities are homogeneous with the same base probability it further simplifies to
$\mathrm{AD}(\nu^{1},\nu^{2})=\mathcal{W}_{*}(\rho^{1},\rho^{2}).$
Example 2.
Let $\tilde{\mathcal{N}}^{1}$ and $\tilde{\mathcal{N}}^{2}$ be two Poisson random measures of intensities $\alpha_{i}P^{i}_{0}$, for $i=1,2$, where $\alpha_{i}>0$ and $P^{i}_{0}$ is a probability on $\mathbb{X}$. Then $\tilde{\mathcal{N}}^{i}$ is a CRM with Lévy intensity $\nu^{i}=\alpha_{i}\delta_{1}(\mathrm{d}s)\mathrm{d}P^{i}_{0}(x)$. In particular they are homogeneous CRMs with $\mathcal{W}_{*}(\rho^{1},\rho^{2})=|\alpha_{1}-\alpha_{2}|$. Thus putting together Theorem 4 and Example 1,
$\mathrm{BL}(\alpha_{1}P^{1}_{0},\alpha_{2}P^{2}_{0})\leq\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mathcal{N}}^{1}),\mathcal{L}(\tilde{\mathcal{N}}^{2}))\leq\frac{\alpha_{1}+\alpha_{2}}{2}\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P^{1}_{0},P^{2}_{0})+|\alpha_{1}-\alpha_{2}|.$
If $P_{0}^{1}=P_{0}^{2}$ then from Example 1 we see that the lower and upper bound coincide, so that
$\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mathcal{N}}^{1}),\mathcal{L}(\tilde{\mathcal{N}}^{2}))=|\alpha_{1}-\alpha_{2}|.$
On the other hand, if $\alpha_{1}=\alpha_{2}=\alpha$, then again from Example 1 we deduce
$\frac{\alpha}{2}\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P^{1}_{0},P^{2}_{0})\leq\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mathcal{N}}^{1}),\mathcal{L}(\tilde{\mathcal{N}}^{2}))\leq\alpha\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P^{1}_{0},P^{2}_{0}),$
which characterizes the value of $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mathcal{N}}^{1}),\mathcal{L}(\tilde{\mathcal{N}}^{2}))$ up to a constant multiplicative factor.
Schuhmacher and Xia (2008) have defined a distance on the laws of point processes with a finite number of points by defining a suitable transport problem tailored for point processes. Interestingly, in Schuhmacher and Xia (2008, Remark 4.2) the authors also provide an upper bound of their distance evaluated on Poisson random measures in terms of the Wasserstein distance between the normalized intensities. Direct comparisons are not meaningful, as their distance and our distance $\mathcal{W}_{\mathrm{BL}}$ differ, but it illustrates the need of relating a distance between Poisson random measures to a distance between their intensities, as the latter is often more tractable and interpretable.
Although $\mathrm{AD}$ is not a distance, because of the presence of $M_{1}(\nu^{1})+M_{1}(\nu^{2})$ in the expression, it becomes a distance if we freeze the value of $M_{1}(\nu)$. Although it may seem like an algebraic trick, in the next section we will see why it is natural to consider the space of Lévy intensities $\nu$ with a given value for $M_{1}(\nu)$.
Proposition 5.
For any $m>0$, $\mathrm{AD}$ defines a distance on the space
$\mathcal{M}_{m}((0,+\infty)\times\mathbb{X})=\left\{\nu\in\mathcal{M}((0,+\infty)\times\mathbb{X})\ :\ M_{1}(\nu)=m\right\}.$
Eventually, for posterior distributions in our Bayesian models, we are led to consider mixtures of CRMs, that is, CRMs with a random Lévy intensity. We call them “Cox CRM”, by analogy with Cox Poisson point processes (Cox, 1955).
Definition 5.
A random measure $\tilde{\mu}$ on $\mathbb{X}$ is a Cox CRM or a doubly stochastic CRM if there exists a random element $\tilde{\nu}$ valued in $\mathcal{M}((0,+\infty)\times\mathbb{X})$ such that $\tilde{\mu}|\tilde{\nu}\sim\textup{CRM}(\tilde{\nu})$.
Note that any $\tilde{\mu}\sim\mathrm{CRM}(\nu)$ is technically a Cox CRM, by taking $\tilde{\nu}$ a.s. constant and equal to $\nu$. Moreover, we observe that $\tilde{\nu}$ is not a random measure according to the definition of Section 3 because it can have infinite mass almost surely. Thanks to the convexity of the Wasserstein distance, it is straightforward to extend the upper bound in Theorem 4.
Theorem 6.
Let $\tilde{\mu}^{i}$ for $i=1,2$ be Cox CRM with random Lévy intensities $\tilde{\nu}^{1}$ and $\tilde{\nu}^{2}$ respectively. If $(\tilde{\nu}^{1},\tilde{\nu}^{2})$ is any coupling between these random Lévy intensities then
$\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}),\mathcal{L}(\tilde{\mu}^{2}))\leq\mathbb{E}(\mathrm{AD}(\tilde{\nu}^{1},\tilde{\nu}^{2})).$
In particular, if we restrict to $\tilde{\nu}^{1},\tilde{\nu}^{2}\in\mathcal{M}_{m}((0,+\infty)\times\mathbb{X})$ for some fixed m¿0, with $\mathcal{W}_{\mathrm{AD}}$ the Wasserstein distance over the $\mathrm{AD}$ metric in Definition 4, we have
$\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}),\mathcal{L}(\tilde{\mu}^{2}))\leq\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1}),\mathcal{L}(\tilde{\nu}^{2})).$ (12)
4 Normalization, identifiability, and scaling
One of the most popular uses of completely random measures is through their normalization $t(\tilde{\mu})=\tilde{\mu}/\tilde{\mu}(\mathbb{X})$, first defined in Regazzini et al. (2003). The distribution of a sequence $(X_{n})_{n\geq 1}$ of exchangeable observations on $\mathbb{X}$ can be flexibly modelled as
$X_{1},\dots,X_{n}|\tilde{\mu}\stackrel{{\scriptstyle\mbox{\scriptsize{\rm iid}}}}{{\sim}}\frac{\tilde{\mu}}{\tilde{\mu}(\mathbb{X})};\qquad\tilde{\mu}\sim\textup{CRM}(\nu),$ (13)
for any $\tilde{\mu}$ that takes values on the space of finite measures $\mathcal{M}_{B}(\mathbb{X})$ and such that $\tilde{\mu}(\mathbb{X})>0$ a.s.. In this case, Regazzini et al. (2003) show that $\tilde{\mu}/\tilde{\mu}(\mathbb{X})$ is a well-defined random probability measure, while James et al. (2006, 2009) find closed-form expressions for the predictive and posterior distribution.
Our goal for the remainder of this work is to use a distance between the laws of random measures $\tilde{\mu}$ to perform comparisons between the induced Bayesian models (13), both a priori and a posteriori. To this end, in this section we investigate which CRMs induce the same normalization in law, a.k.a., the identifiability of the law of a CRM. Theorem 7 finds all equivalent CRMs, in the sense that they induce the same normalization. The choice of a representative for each equivalence class leads us to the notion of scaled CRM. The analysis is then extended to Cox CRMs, which naturally emerge in the study of the posterior, as will be clear in Section 5.
Remark 3.
For the normalization $\tilde{\mu}/\tilde{\mu}(\mathbb{X})$ to be well-defined one needs $\tilde{\mu}(\mathbb{X})>0$ a.s., which at the level of the Lévy intensities corresponds to $\nu((0,+\infty)\times\mathbb{X})=+\infty$ (Regazzini et al., 2003). We will require a slightly stronger condition, namely that the CRM is infintely active. In terms of the canonical decomposition of Proposition 3, this assumption reads
for $\text{for }P_{0}\text{-a.e. }x,\qquad\rho_{x}(0,+\infty)=+\infty.$
This means that there will always be an infinite number of jumps in a set $A$ with $P_{0}(A)>0$, and thus the compound Poisson approximation (9) cannot be extended to $\epsilon=0$. In particular, it guarantees that for any set $A$ exactly one of the two alternatives holds: either $P_{0}(A)>0$, and in this case $\tilde{\mu}(A)>0$ a.s., or $P_{0}(A)=0$, which implies $\tilde{\mu}(A)=0$ a.s. This will be useful for our identifiability results.
4.1 Completely random measures
Quite clearly, different CRMs $\tilde{\mu}$ can lead to the same normalized random measure $\tilde{\mu}/\tilde{\mu}(\mathbb{X})$: for example, $\tilde{\mu}$ and $\alpha\tilde{\mu}$ lead to the same normalization for any $\alpha>0$. In Theorem 7 below we build on a result on subordinators by David Aldous and Stevan Evans (Pitman and Yor, 1992, Lemma 7.5) to prove that this can be the only source for the lack of identifiability.
Theorem 7.
Let $\tilde{\mu}^{i}\sim\textup{CRM}(\nu^{i})$ be infinitely active with finite mean, for $i=1,2$. Assume that $\tilde{\mu}^{i}$ is non-degenerate, in the sense $P^{i}_{0}$ is not a Dirac mass, for $i=1,2$. Then,
$\frac{\tilde{\mu}^{1}}{\tilde{\mu}^{1}(\mathbb{X})}\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\frac{\tilde{\mu}^{2}}{\tilde{\mu}^{2}(\mathbb{X})}\qquad\text{ if and only if }\qquad\tilde{\mu}^{1}\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\alpha\tilde{\mu}^{2}\text{ for some scalar }\alpha>0.$
The assumption of non-degeneracy is only here to exclude the case $\tilde{\mu}^{i}=J_{i}\delta_{x}$ for $i=1,2$ with $x$ deterministic and $J_{1},J_{2}$ random. In fact, in this case $\tilde{\mu}^{1}/\tilde{\mu}^{1}(\mathbb{X})=\tilde{\mu}^{2}/\tilde{\mu}^{2}(\mathbb{X})=\delta_{x}$ a.s., even though it does not hold $\tilde{\mu}^{1}\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\alpha\tilde{\mu}^{2}$ for some $\alpha>0$. Theorem 7 sheds new light on the equivalence relation $\sim$ that identifies CRMs leading to the same normalization, showing that $\tilde{\mu}^{1}\sim\tilde{\mu}^{2}$ if and only if $\tilde{\mu}^{1}\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\alpha\tilde{\mu}^{2}$ for some $\alpha>0$. We use it to choose one representative for each equivalence class, namely the only CRM in the class with total expected value equal to 1, as in the next definition.
Definition 6.
A CRM $\tilde{\mu}$ is a scaled CRM if $\mathbb{E}(\tilde{\mu}(\mathbb{X}))=1$. If $\tilde{\mu}$ is a CRM with finite mean, we define its scaled CRM as
$\tilde{\mu}_{\mathcal{S}}=\frac{\tilde{\mu}}{\mathbb{E}(\tilde{\mu}(\mathbb{X}))}.$
Corollary 8.
Let $\tilde{\mu}^{i}\sim\textup{CRM}(\nu^{i})$ be infinitely active with finite mean, for $i=1,2$. Then,
$\frac{\tilde{\mu}^{1}}{\tilde{\mu}^{1}(\mathbb{X})}\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\frac{\tilde{\mu}^{2}}{\tilde{\mu}^{2}(\mathbb{X})}\qquad\text{if and only if}\qquad\tilde{\mu}_{\mathcal{S}}^{1}\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\tilde{\mu}_{\mathcal{S}}^{2}.$
By Proposition 3 the mean measure $Q(\cdot)=\mathbb{E}(\tilde{\mu}(\cdot))$ of any scaled CRM with Lévy intensity $\mathrm{d}\nu(s,x)=\mathrm{d}\rho_{x}(s)\mathrm{d}P_{0}(x)$ is the probability measure $P_{0}$ and, for $P_{0}$-a.e. $x\in\mathbb{X}$,
$\int_{0}^{+\infty}s\,\mathrm{d}\rho_{x}(s)=1.$ (14)
We highlight that, unlike normalized CRMs, scaled CRMs are CRMs. The following result provides a general way to find their Lévy intensities. We state it here for diffuse jumps components $\mathrm{d}\rho_{x}(s)=\rho_{x}(s)\,\mathrm{d}s$ for simplicity, though it can be extended to atomic jump components; see Lemma SM3.
Lemma 9.
Let $\tilde{\mu}\sim\textup{CRM}(\nu)$ such that $\mathrm{d}\nu(s,x)=\rho_{x}(s)\,\mathrm{d}s\,\mathrm{d}P_{0}(x)$. Then $\tilde{\mu}_{\mathcal{S}}$ has Lévy intensity
$\mathrm{d}\nu_{\mathcal{S}}(s,x)=\mathbb{E}(\tilde{\mu}(\mathbb{X}))\rho_{x}(\mathbb{E}(\tilde{\mu}(\mathbb{X}))s)\,\mathrm{d}s\,\mathrm{d}P_{0}(x).$
Example 3 (Gamma CRM).
A gamma CRM with base probability $P_{0}\in\mathcal{P}(\mathbb{X})$, total base measure $\alpha>0$ and scale parameter $b>0$ has Lévy intensity $\mathrm{d}\nu(s,x)=\alpha e^{-bs}s^{-1}\mathrm{d}s\,\mathrm{d}P_{0}(x)$. Thanks to Campbell’s theorem one easily shows that $\mathbb{E}(\tilde{\mu}(\mathbb{X}))=\alpha/b$, so that by Lemma 9 the corresponding scaled gamma CRM has Lévy intensity
$\mathrm{d}\nu_{\mathcal{S}}(s,x)=\alpha\frac{e^{-\alpha s}}{s}\mathrm{d}s\,\mathrm{d}P_{0}(x).$ (15)
Thus, $\tilde{\mu}_{\mathcal{S}}$ is a gamma CRM with base probability $P_{0}$, total base measure $\alpha$ and scale parameter $\alpha$. This highlights the redundancy of the parameter $b>0$, which will be assumed to be equal to 1 in the remainder of the paper, unless otherwise specified.
Example 4 (Generalized gamma CRM).
Another popular prior in BNP models is the normalized generalized gamma (Lijoi et al., 2007, Barrios et al., 2013). The generalized gamma CRM with total base measure $\alpha>0$, discount parameter $\sigma\in(0,1)$, and base probability $P_{0}\in\mathcal{P}(\mathbb{X})$ has Lévy intensity
$\mathrm{d}\nu(s,x)=\frac{\alpha}{\Gamma(1-\sigma)}\frac{e^{-s}}{s^{1+\sigma}}\mathrm{d}s\,\mathrm{d}P_{0}(x),$
where $\Gamma(z)=\int_{0}^{+\infty}x^{z-1}e^{-x}\,\mathrm{d}x$ is the gamma function. In particular, the gamma CRM is recovered when $\sigma=0$ and the inverse Gaussian CRM (Lijoi et al., 2005) when $\sigma=1/2$. It is easy to prove that $\mathbb{E}(\tilde{\mu}(\mathbb{X}))=\alpha$, so that by Lemma 9 the corresponding scaled CRM has Lévy intensity
$\mathrm{d}\nu_{\mathcal{S}}(s,x)=\frac{\alpha^{1-\sigma}}{\Gamma(1-\sigma)}\frac{e^{-\alpha s}}{s^{1+\sigma}}\mathrm{d}s\,\mathrm{d}P_{0}(x).$ (16)
Thus we can regard $\tilde{\mu}_{\mathcal{S}}$ as a generalized gamma CRM that also accounts for a scale parameter.
When comparing normalized CRMs we should thus restrict their scaled version. Importantly, Proposition 5 and Theorem 4 guarantee that
$\mathrm{AD}(\nu^{1}_{\mathcal{S}},\nu^{2}_{\mathcal{S}})=\inf_{\pi\in\Pi(P_{0}^{1},P_{0}^{2})}\;\mathbb{E}_{(X,Y)\sim\pi}\left(d_{\mathbb{X},\text{t}}(X,Y)+\mathcal{W}_{*}(\rho_{X}^{1},\rho_{Y}^{2})\right)$ (17)
is a distance on on Lévy intensities of scaled CRMs that satisfies $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2}_{\mathcal{S}}))\leq\mathrm{AD}(\nu^{1}_{\mathcal{S}},\nu^{2}_{\mathcal{S}})$. Since Lévy intensities uniquely characterize the law of the CRM, the distance $\mathrm{AD}$ can be considered as a distance between CRMs: arguably a less canonical but more tractable one.
4.2 Cox completely random measures
We now extend the notion of scaling and identifiability to Cox completely random measures.
Definition 7.
Let $\tilde{\mu}$ be a Cox CRM such that $\tilde{\mu}|\tilde{\nu}\sim\textup{CRM}(\tilde{\nu})$ has a.s. finite mean and infinite activity. Then we define its Cox scaled CRM $\tilde{\mu}_{\mathcal{S}}$ as
$\tilde{\mu}_{\mathcal{S}}|\tilde{\nu}=\frac{\tilde{\mu}}{\mathbb{E}(\tilde{\mu}(\mathbb{X})|\tilde{\nu})}.$
Theorem 10.
Let $\tilde{\mu}^{i}$ be a Cox CRM such that $\tilde{\mu}^{i}|\tilde{\nu}^{i}\sim\textup{CRM}(\tilde{\nu}^{i})$ is a.s. an infinitely active CRM with finite mean, for $i=1,2$. Denoting $\mathrm{d}\tilde{\nu}^{i}(s,x)=\mathrm{d}\tilde{\rho}^{i}(s)\,\mathrm{d}\tilde{P}^{i}_{0}(x)$, we assume that i) $\mathbb{E}(\tilde{P}^{i}_{0})$ is not purely atomic, (ii) for any Borel set $A$, either $\tilde{P}^{i}_{0}(A)>0$ a.s. or $\tilde{P}^{i}_{0}(A)=0$ a.s.. Then,
$\frac{\tilde{\mu}^{1}}{\tilde{\mu}^{1}(\mathbb{X})}\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\frac{\tilde{\mu}^{2}}{\tilde{\mu}^{2}(\mathbb{X})}\qquad\text{if and only if}\qquad\tilde{\mu}_{\mathcal{S}}^{1}\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\tilde{\mu}_{\mathcal{S}}^{2}.$
Following Proposition 3, we see that if $\tilde{\mu}|\tilde{\nu}\sim\mathrm{CRM}(\tilde{\nu})$, then it is a Cox scaled CRM if $\tilde{\nu}\in\mathcal{M}_{1}((0,+\infty)\times\mathbb{X})$ a.s. We denote by $\tilde{\nu}_{\mathcal{S}}$ the random Lévy measure of the Cox scaled CRM, also referred to as random scaled Lévy measure.
Remark 4.
Theorem 10 is the analogue for Cox CRMs of Corollary 8 for CRMs. However, it is worth noting that the latter does not follow directly from Theorem 10. Indeed, when dealing with CRMs we do not need the extra assumption that $P_{0}$ is not purely atomic.
By reasoning as for CRMs, thanks to Theorem 10 when comparing normalized Cox CRMs we should restrict their scaled version. Theorem 6 ensures that $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2}_{\mathcal{S}}))$ is a distance on random scaled Lévy measures that satisfies $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2}_{\mathcal{S}}))\leq\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2}_{\mathcal{S}}))$.
5 Merging rate of Bayesian nonparametric models
In this section we use our distance to measure the impact of the prior for Bayesian nonparametric models of the form (13). Specifically, we consider the distance between the posteriors corresponding to two different priors and (i) we determine if the distance vanishes as the number of observations increases, (ii) if it vanishes, we analyze its rate of convergence to 0. This is what we call the merging rate of opinions. In Section 5.1 we use this general framework to study the impact of the parameters of the Dirichlet processes, whereas in Section 5.2 we investigate the impact of the discount parameter $\sigma$ of the generalized gamma CRM, showing the exceptional differences between these two scenarios.
The remarkable Theorem 1 in James et al. (2009) provides a regular version of the posterior $\mathcal{L}(\tilde{\mu}|X_{1:n})$ arising from model (13), that is, a Markov kernel $\phi:\mathbb{X}^{n}\times\mathcal{B}(\mathcal{M})\to[0,1]$, where $\mathcal{B}(\mathcal{M})$ denotes the Borel $\sigma$-algebra on the space of locally finite measures $\mathcal{M}(\mathbb{X})$. For observed data $x_{1:n}$, we denote by $\mathcal{L}(\tilde{\mu}|X_{1:n}=x_{1:n})$ the probability on $\mathcal{M}(\mathbb{X})$ defined by $A\mapsto\phi(x_{1:n},A)$. A crucial aspect of our setting is that $x_{1:n}$ is a generic deterministic sequence. It could be a realization from the model, a realization from another model, or indeed any deterministic sequence of data: we are able to analyze the asymptotic behaviour of the merging without putting any distributional assumptions on $x_{1:n}$. We will explain how our results depend only on the number of observations $n$ and the number $k=k_{n}$ of distinct values among the $n$ first observations.
We denote by $\tilde{\mu}^{*}\sim\mathcal{L}(\tilde{\mu}|X_{1:n}=x_{1:n})$ a random measure that we often address as the posterior random measure, omitting the dependence on $n$ for notational convenience. The aforementioned Theorem 1 in James et al. (2009), which we state below for completeness, shows that $\tilde{\mu}^{*}$ is a CRM conditionally on a latent variable. Using the terminology of Section 3, this implies that $\tilde{\mu}^{*}$ is a Cox CRM. To preserve the identifiability discussed in Section 4, we do not define the distance at the level of the posterior Cox CRM $\tilde{\mu}^{*}$ but rather on its scaled version $\tilde{\mu}^{i*}_{\mathcal{S}}$ (cf. Definition 7). We focus on the distance $\mathcal{W}_{\mathrm{AD}}$ between their random Lévy measures, where $\mathcal{W}_{\mathrm{AD}}$ is the Wasserstein distance over the $\mathrm{AD}$ metric in Definition 4, which is tailored for Cox scaled CRMs and whose explicit expressions will allow us to derive exact asymptotic behaviours. As described in (12), these can be used to derive upper bounds of $\mathcal{W}_{\mathrm{BL}}$, the Wasserstein distance over the $\mathrm{BL}$ metric in (5), which is a natural distance for arbitrary random measures without additional structure.
In the rest of the section the base probability $P_{0}$ is assumed to be non-atomic, $(x_{1}^{*},\dots,x_{k}^{*})$ are the $k\leq n$ distinct observations in $(x_{1},\dots,x_{n})$ and $n_{i}=\#\{j:x_{j}=x_{i}^{*}\}$ is the number of observations equal to $x_{i}^{*}$. We define a random variable $U$ with the following p.d.f.
$f_{U}(u)\propto u^{n-1}e^{-\psi(u)}\prod_{i=1}^{k}\tau_{n_{i}|x_{i}^{*}}(u),$ (18)
where $\psi$ is the Laplace exponent of $\tilde{\mu}$ and $\tau_{m|x}$ are its cumulants, respectively equal to
$\psi(u)=\iint_{(0,+\infty)\times\mathbb{X}}(1-e^{-us})\,\mathrm{d}\nu(s,x),\qquad\tau_{m|x}(u)=\int_{(0,+\infty)}e^{-us}s^{m}\mathrm{d}\rho_{x}(s).$
Theorem 11 (James et al. (2009)).
Let $\tilde{\mu}$ be a CRM with Lévy intensity $\mathrm{d}\nu(s,x)=\mathrm{d}\rho_{x}(s)\mathrm{d}P_{0}(x)$. Consider $U=U_{n}$ a latent variable with p.d.f. (18) and define $\tilde{\mu}^{*}$ such that
$\tilde{\mu}^{*}|U\stackrel{{\scriptstyle\mbox{\scriptsize{{d}}}}}{{=}}\tilde{\mu}_{U}+\sum_{i=1}^{k}J_{i}^{U}\delta_{x_{i}^{*}},$
where (i) $\tilde{\mu}_{U}$ is a CRM with Lévy intensity $\mathrm{d}\nu_{U}(s,x)=\mathrm{e}^{-Us}\mathrm{d}\rho_{x}(s)\mathrm{d}P_{0}(x)$, (ii) the i-th jump $J_{i}^{U}$ has p.d.f. $f_{i}(s)\propto$ $s^{n_{i}}\mathrm{e}^{-Us}\mathrm{d}\rho_{x_{i}^{*}}(s)$, (iii) $\{J_{i}^{U}\}_{i}$ are mutually independent and independent from $\tilde{\mu}_{U}$.
Then $\mathcal{L}(\tilde{\mu}^{*})=\mathcal{L}(\tilde{\mu}|X_{1:n}=x_{1:n})$ is a version of the posterior distribution according to model (13).
5.1 Dirichlet process
In this section we study the impact of the parameters of the Dirichlet process, which is a normalized gamma CRM, by studying the OT distance between scaled posteriors. We first show that the posterior is a Cox CRM and find its corresponding scaled version (Lemma 12). Then, we prove several finite-sample and asymptotic properties for the OT distance (Proposition 14). This allows us to state new conclusions about the merging rate of opinions (Corollary 15) and study its finite-sample behaviour.
First of all we show that the posterior is a Cox CRM. This result builds on Theorem 11, by further showing that the jumps $J_{i}^{U}$ are infinitely divisible, and can thus be included in the random Lévy intensity, as shown in Lemma SM4 in the supplementary material.
Lemma 12.
Let $\tilde{\mu}$ be a gamma CRM of parameters $(\alpha,P_{0})$. Then $\tilde{\mu}^{*}$ is a Cox CRM and $\tilde{\mu}^{*}_{\mathcal{S}}$ is a scaled gamma CRM with canonical Lévy intensity $\mathrm{d}\nu^{*}_{\mathcal{S}}(s,x)=\rho^{*}(s)\mathrm{d}s\,\mathrm{d}P_{0}^{*}(x)$, where
$\rho^{*}(s)=(\alpha+n)\frac{e^{-(\alpha+n)s}}{s};\qquad P_{0}^{*}=\frac{\alpha}{\alpha+n}P_{0}+\frac{1}{\alpha+n}\sum_{i=1}^{n}\delta_{x_{i}}.$
Lemma 12 shows that the posterior of a gamma CRM is a Cox CRM with a peculiar feature: all the randomness of its Lévy intensity is contained in a scale parameter, so that the corresponding Cox scaled CRM reduces to a scaled CRM. In the sequel we write $\Gamma(z,t)=\int_{t}^{+\infty}x^{1-z}e^{-x}\,\mathrm{d}x$ for the incomplete Gamma function.
Proposition 13.
Let $\tilde{\mu}^{i}$ be a gamma CRM of parameters $(\alpha_{i},P_{0}^{i})$ with $P_{0}^{i}$ having finite first moment for $i=1,2$. Then, $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2}_{\mathcal{S}}))=\mathrm{AD}(\nu^{1*}_{\mathcal{S}},\nu^{2*}_{\mathcal{S}})=\mathcal{J}+\mathcal{A}$, with $\mathcal{J}=\mathcal{J}(\alpha_{1},\alpha_{2},n)$ and $\mathcal{A}=\mathcal{A}(\alpha_{1},P_{0}^{1},\alpha_{2},P_{0}^{2},n)$ defined as
$\mathcal{J}$ $=\int_{0}^{+\infty}|(\alpha_{1}+n)\Gamma(0,(\alpha_{1}+n)s)-(\alpha_{2}+n)\Gamma(0,(\alpha_{2}+n)s)|\mathrm{d}s;$
$\mathcal{A}$ $=\mathcal{W}_{d_{\mathbb{X},\text{t}}}\bigg{(}\frac{\alpha_{1}}{\alpha_{1}+n}P_{0}^{1}+\frac{1}{\alpha_{1}+n}\sum_{i=1}^{n}\delta_{x_{i}},\frac{\alpha_{2}}{\alpha_{2}+n}P_{0}^{2}+\frac{1}{\alpha_{2}+n}\sum_{i=1}^{n}\delta_{x_{i}}\bigg{)}.$
We observe that $\mathcal{J}$ measures the discrepancy between the jumps of the posterior random measures, which does not depend on the base probability measures and depends on the observations only through their cardinality $n$. As for $\mathcal{A}$, it measures the discrepancy between the atoms a posteriori and it coincides with the truncated Wasserstein distance between the predictive distributions $\mathcal{L}(X_{n+1}|X_{1:n})$ of the corresponding models (13).
Proposition 14.
Consider the framework of Proposition 13 where, without loss of generality, $\alpha_{1}\leq\alpha_{2}$, and let $C=\int_{0}^{+\infty}|\Gamma(0,t)-e^{-t}|\mathrm{d}t$. Then, there holds
  1. 1.
    $n\mapsto\mathcal{J}(\alpha_{1},\alpha_{2},n)$ is decreasing in $n$.
  2. 2.
    For every $n\geq 0$,
    $\mathcal{J}(\alpha_{1},\alpha_{2},n)\leq C\log\bigg{(}\frac{\alpha_{2}+n}{\alpha_{1}+n}\bigg{)}.$ (19)
  3. 3.
    As $n$ goes to +$\infty$+ ∞,
    $\mathcal{J}(\alpha_{1},\alpha_{2},n)=\frac{C(\alpha_{2}-\alpha_{1})}{n}+o\bigg{(}\frac{1}{n}\bigg{)}.$
  4. 4.
    For every $n\geq 1$, the following inequality holds, with equality if $P_{0}^{1}=P_{0}^{2}$ or $\alpha_{1}=\alpha_{2}$,
    $\mathcal{A}\leq\frac{\alpha_{1}}{\alpha_{1}+n}\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0}^{1},P_{0}^{2})+\frac{(\alpha_{2}-\alpha_{1})n}{(\alpha_{1}+n)(\alpha_{2}+n)}\mathcal{W}_{d_{\mathbb{X},\text{t}}}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}},P_{0}^{2}\bigg{)}.$
Proposition 14 gives us the means to unravel both finite-sample and asymptotic properties of the distance between posterior random measures. From an asymptotic point of view, the next result is an immediate but remarkable consequence. For sequences $(a_{n})_{n\in\mathbb{N}}$, $(b_{n})_{n\in\mathbb{N}}$ we write $a_{n}\asymp b_{n}$ when the quotient $(a_{n}/b_{n})_{n\in\mathbb{N}}$ stays bounded from above and from below by a strictly positive constant.
Theorem 15.
Let $\tilde{\mu}^{1},\tilde{\mu}^{1}$ be two gamma CRMs of parameters $(\alpha_{i},P_{0}^{i})$, for $i=1,2$, with $\alpha_{1}\neq\alpha_{2}$ or $P_{0}^{1}\neq P_{0}^{2}$. Then for any sequence $(x_{n})_{n\geq 1}$ of observations on $\mathbb{X}$,
$\mathrm{AD}(\nu^{1*}_{\mathcal{S}},\nu^{2*}_{\mathcal{S}})\asymp\frac{1}{n}.$
In particular, $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2*}_{\mathcal{S}}))\lesssim 1/n$.
Theorem 15 states that there will always be merging of opinions, and provides us with the merging rate of $1/n$, which holds independently of the choice of parameters $(\alpha_{i},P_{0}^{i})$ and independently of the dimension of the space $\mathbb{X}$. Moreover, thanks to Proposition 14 we can also study the starting time of the merging, which highlights a different behaviour for the jump and the atom component, as discussed in the remainder of this section.
Figure 1: Jump component $\mathcal{J}$ of $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))$ when $\tilde{\mu}^{1}$ is a gamma$(\alpha_{1}=1,P_{0}^{1})$ CRM and $\tilde{\mu}^{2}$ is a gamma$(\alpha_{2},P_{0}^{2})$ CRM, as the number of observations $n$ increases. The upper bounds are the right hand side of (19).
Point 1. of Proposition 14 states that the jump component $\mathcal{J}$ decreases steadily in $n$, that is, the merging starts with the first observation (time 1). We illustrate this behavior in Figure 1, where without loss of generality we fix $\alpha=1$ and plot the value of $\mathcal{J}$ as a function of $n$, for different values of $\alpha_{2}$, together with the upper bound (19), which becomes asymptotically exact as $n\to+\infty$, as stated in point 3. of Proposition 14. As for the atom component $\mathcal{A}$, this can show different behaviours depending on the parameters $(\alpha_{i},P_{0}^{i})$, as showcased by the study of the upper bounds in point 4. of Proposition 14. In particular, if $\alpha_{1}=\alpha_{2}=\alpha$,
$\mathcal{A}=\frac{\alpha}{\alpha+n}\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0}^{1},P_{0}^{2}).$
Thus in this scenario $\mathcal{A}$ steadily decreases as $n$ increases, that is, the merging of the atoms starts at time 1. The left panel of Figure 2 validates it numerically by taking $\alpha_{1}=\alpha_{2}=\alpha$ for different values of $\alpha$, $P_{0}^{1}$ and $P_{0}^{2}$ Gaussian distributions with different mean and same variance, and i.i.d. observations from a Poisson distribution.
Figure 2: Atom component $\mathcal{A}$ of $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))$, as the number of observations $n$ increases, in two different scenarios. Left: $\tilde{\mu}^{1}$ is a gamma$(\alpha_{1}=\alpha,P_{0}^{1}=\mathcal{N}(1,1))$ CRM and $\tilde{\mu}^{2}$ is a gamma$(\alpha_{2}=\alpha,P_{0}^{2}=\mathcal{N}(2,1))$ CRM, where $\mathcal{N}(m,v)$ is a Gaussian distribution of mean $m$ and variance $v$. Right: $\tilde{\mu}^{1}$ is a gamma$(\alpha_{1}=10,P_{0}^{1}=\mathcal{N}(1,1))$ CRM and $\tilde{\mu}^{2}$ is a gamma$(\alpha_{2},P_{0}^{2}=\mathcal{N}(1,1))$ CRM. In both scenarios the observations are an i.i.d. sample from a Poisson distribution of mean 1.
If on the other hand $P^{1}_{0}=P^{2}_{0}$,
$\mathcal{A}=\frac{(\alpha_{2}-\alpha_{1})n}{(\alpha_{1}+n)(\alpha_{2}+n)}\mathcal{W}_{d_{\mathbb{X},\text{t}}}\bigg{(}P_{0},\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}\bigg{)}.$
For many standard data-generating mechanisms for the data, i.e., if we consider $(x_{n})_{n\geq 1}$ as a realization of a random sequence $(X_{n})_{n\geq 1}$ possibly differing from model (13), $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}})$ has a non-zero limit almost surely. For example, this happens if the data are i.i.d. samples from $P\neq P_{0}$ thanks to the law of large numbers or if they are exchangeable thanks to de Finetti Theorem. In such case, for $n$ large enough, the behaviour of $\mathcal{A}$ with respect to $n$ reduces to the behaviour of the prefactor
$\omega(n)=\frac{(\alpha_{2}-\alpha_{1})n}{(\alpha_{1}+n)(\alpha_{2}+n)}.$
This quantity can be studied analytically: it first increases and reaches a maximum at $n\approx\sqrt{\alpha_{1}\alpha_{2}}$ before decreasing to 00 at rate $1/n$, hinting that this may well be the case for $\mathcal{A}$ as well, at least when $\sqrt{\alpha_{1}\alpha_{2}}$ is large enough. We tested this behavior numerically. In Figure 2 on the right we fix $P_{0}^{1}=P_{0}^{2}=P_{0}$ to be a Gaussian distribution, $\alpha_{1}=10$ and $\alpha_{2}$ to vary, while the observations are i.i.d. from $P$ a Poisson distribution: in this setting we know that the expected value of $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}})$ converges to $\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},P)$ with rate $1/\sqrt{n}$ (Fournier and Guillin, 2015). We see that the distance is not decreasing in $n$, the maximum being achieved for $\bar{n}\approx 34$ when $\alpha_{2}=50$, $\bar{n}\approx 42$ when $\alpha_{2}=100$ and $\bar{n}\approx 64$ when $\alpha_{2}=500$. As expected, the maximum is better approximated by $\sqrt{\alpha_{1}\alpha_{2}}$ (equal to approximately 21, 31, and 72, respectively) as $\sqrt{\alpha_{1}\alpha_{2}}$ increases.
In conclusion, depending on the total base measure and the base probabilities, we can have different behaviours: either a steady decrease, or an increase phase followed by a decrease, but ultimately the asymptotic rate of convergence is $1/n$. From a statistical perspective this finding is quite remarkable: even if two Bayesians have the same prior on a phenomenon, the difference between their opinions may increase if they see a finite amount of data, according to how much confidence they are willing to put on their prior guesses.
5.2 Normalized generalized gamma process
In this section we study the impact of the discount parameter $\sigma$ of a generalized gamma CRM, that is, the effect of using a normalized generalized gamma CRM instead of a Dirichlet process in model (13). As in the case of Gamma CRMs, we first prove that the posterior is a Cox CRM and find the expression of its Cox scaled CRM (Proposition 16), which is used to express the OT distance between the posteriors in Corollary 17. We then develop an in-depth asymptotic study of the OT distance as the number of observations diverge. Since the Lévy intensities depend on a latent variable $U=U_{n}$, we first need to unravel its asymptotic behaviour, which crucially depends on the number of distinct observations $k=k_{n}$ (Theorem 18). This gives us the means to study its implications on the behaviour on the merging of opinions (Theorem 19 and Proposition 20).
Recall that $(x_{1}^{*},\dots,x_{k}^{*})$ are the $k\leq n$ distinct observations in $(x_{1},\dots,x_{n})$. We show that the posterior is a Cox CRM with the same technique used in Section 5.1 for the gamma CRM. We define
$c=c(\alpha,\sigma,n,k,U)=\alpha(1+U)^{\sigma}+n-k\sigma.$
Proposition 16.
Let $\tilde{\mu}$ be a generalized gamma CRM of parameters $(\alpha,\sigma,P_{0})$. Then $\tilde{\mu}^{*}$ is a Cox CRM and $\tilde{\mu}^{*}_{\mathcal{S}}$ is a Cox scaled CRM whose random Lévy intensity has canonical form $\mathrm{d}\tilde{\nu}_{\mathcal{S}}^{*}(s,x)=\tilde{\rho}^{*}_{x}(s)\mathrm{d}s\,\mathrm{d}\tilde{P}_{0}^{*}(x)$, where
$\tilde{\rho}^{*}_{x}(s)=\frac{c^{1-\sigma}}{\Gamma(1-\sigma)}\frac{e^{-cs}}{s^{1+\sigma}}\mathbbm{1}_{\mathbb{X}\setminus\{x_{1}^{*},\dots,x_{k}^{*}\}}(x)+c\frac{e^{-cs}}{s}\mathbbm{1}_{\{x_{1}^{*},\dots,x_{k}^{*}\}}(x),$
$\tilde{P}_{0}^{*}=\frac{\alpha(1+U)^{\sigma}}{c}P_{0}+\frac{1}{c}\sum_{i=1}^{k}(n_{i}-\sigma)\delta_{x_{i}^{*}},$
and $U=U_{n}$ has probability density proportional to $u^{n-1}(1+u)^{k\sigma-n}e^{-\frac{\alpha}{\sigma}(1+u)^{\sigma}}\mathbbm{1}_{(0,+\infty)}(u)$.
We then use this expression to compute the OT distance a posteriori between a generalized gamma and a gamma CRM.
Proposition 17.
Let $\tilde{\mu}^{1}$ be a generalized gamma CRM of parameters $(\alpha,\sigma,P_{0})$, and let $\tilde{\mu}^{2}$ be a gamma CRM with same total base measure $\alpha$ and base probability $P_{0}$. Then,
$\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))=\mathbb{E}_{U}(\mathcal{J}_{1}^{U}+\mathcal{J}_{2}^{U}+\mathcal{A}^{U}),$
where
$\mathcal{J}_{1}^{U}=\frac{\alpha(1+U)^{\sigma}}{c}\mathcal{W}_{*}\bigg{(}\frac{c^{1-\sigma}}{\Gamma(1-\sigma)}\frac{e^{-cs}}{s^{1+\sigma}}\mathrm{d}s,(\alpha+n)\frac{e^{-(\alpha+n)s}}{s}\mathrm{d}s\bigg{)},$
$\mathcal{J}_{2}^{U}=\frac{n-k\sigma}{c}\mathcal{W}_{*}\bigg{(}c\frac{e^{-cs}}{s}\mathrm{d}s,(\alpha+n)\frac{e^{-(\alpha+n)s}}{s}\mathrm{d}s\bigg{)},$
$\mathcal{A}^{U}=\mathcal{W}_{d_{\mathbb{X},\text{t}}}\bigg{(}\frac{\alpha(U+1)^{\sigma}}{c}P_{0}+\frac{1}{c}\sum_{i=1}^{k}(n_{i}-\sigma)\delta_{X_{i}^{*}},\frac{\alpha}{\alpha+n}P_{0}+\frac{1}{\alpha+n}\sum_{i=1}^{k}n_{i}\delta_{x_{i}^{*}}\bigg{)},$
with $U=U_{n}$ and $c$ are defined as in Proposition 16.
Remark 5.
The expression in Proposition 17 can be approximated through Monte Carlo integration, that is, by taking $u_{1},\dots,u_{N}\stackrel{{\scriptstyle\mbox{\scriptsize{\rm iid}}}}{{\sim}}\mathcal{L}(U)$ and then evaluating the unbiased estimator
$\widehat{\mathcal{W}_{\mathrm{AD}}}=\frac{1}{N}\sum_{i=1}^{N}(\mathcal{J}_{1}^{u_{i}}+\mathcal{J}_{2}^{u_{i}}+\mathcal{A}^{u_{i}}),$
where $\mathcal{J}_{1}^{u_{i}}$ and $\mathcal{J}_{2}^{u_{i}}$ can be evaluated by numerical integration and $\mathcal{A}^{u_{i}}$ only involves the classical 1-Wasserstein distance on $\mathbb{X}$ with a truncated metric, which can be approximated by many existing efficient algorithms (Peyré and Cuturi, 2019).
To investigate general features of $\mathcal{W}_{\mathrm{AD}}$ as the sample size $n$ grows, one needs to study the asymptotic behaviour of the latent variables $U=U_{n}$. The next result finds the asymptotic distribution of $(1+U)^{\sigma}$ as the number of observations diverge. We focus on $(1+U)^{\sigma}$ instead of $U$ because the former is featured in the expressions of Proposition 17. The core of the proof shows that $(1+U)^{\sigma}$ has a log-concave density, and concentrates around a point of maximum of its log-density. Indeed, with a change of variable one finds that the density of $(1+U)^{\sigma}$ is proportional to $\exp(-f_{n}(x))$, where
$f_{n}(x)=-(k_{n}-1)\log(x)+\frac{\alpha}{\sigma}x-(n-1)\log\bigg{(}1-\frac{1}{x^{1/\sigma}}\bigg{)}$
is a convex function. If $f_{n}$ were proportional to $n$, say $f_{n}(x)=ng(x)$ for some smooth function $g$, then the result would come directly from Laplace’s method. Here the situation is more delicate as there are different scales (namely, the constant $\alpha/\sigma$, $k_{n}$ and $n$) in the function $f_{n}$. We extend Laplace’s method to this setting in the supplementary material, which yields a general result of independent interest. Note on the other hand that the density of $U$ itself is not log-concave.
Before stating the result let us first introduce the following standard notations for positive deterministic sequences $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$:
  • $a_{n}\ll b_{n}$ if $(a_{n}/b_{n})_{n\in\mathbb{N}}$ converges to 00. Equivalently, $a_{n}=o(b_{n})$.
  • $a_{n}\lesssim b_{n}$ if $(a_{n}/b_{n})_{n\in\mathbb{N}}$ stays bounded from above.
  • $a_{n}\sim b_{n}$ if $(a_{n}/b_{n})_{n\in\mathbb{N}}$ converges to $1$.
  • $a_{n}\asymp b_{n}$ if $(a_{n}/b_{n})_{n\in\mathbb{N}}$ stays bounded from above and from below by a strictly positive constant. Equivalently, $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
Theorem 18.
As $n\rightarrow+\infty$, for the $L^{1}$ convergence,
$\lim_{n\to+\infty}\frac{(1+U)^{\sigma}}{r_{n}}=1,$
where the sequence $(r_{n})_{n\in\mathbb{N}}$ is defined by
$r_{n}=\begin{cases}\alpha^{-\frac{\sigma}{1+\sigma}}\ n^{\frac{\sigma}{1+\sigma}}&\text{\qquad if }k\ll n^{\frac{\sigma}{1+\sigma}},\\ \gamma\ n^{\frac{\sigma}{1+\sigma}}&\text{\qquad if }k\sim\lambda n^{\frac{\sigma}{1+\sigma}},\\ \sigma/\alpha\ k&\text{\qquad if }k\gg n^{\frac{\sigma}{1+\sigma}},\end{cases}$
being here $\gamma$ is the unique solution of $\sigma\lambda x^{-1}+x^{-(1+\sigma)/\sigma}=\alpha$.
Theorem 18 highlights a transition in the asymptotic behaviour of $(1+U)^{\sigma}$ as the number of unique values $k$ increases. Indeed, one can easily derive that
$\mathbb{E}((1+U)^{\sigma})\asymp\max\left(n^{\frac{\sigma}{1+\sigma}},k\right).$
This implies that the growth rate of $(1+U)^{\sigma}$ is not affected by $k$ if $k$ does not diverge or diverge sufficiently slow, otherwise it coincides with $k$ itself, as shown in Figure 3.
We now use Theorem 18 to study the asymptotic behaviour of the distance in Proposition 17.
Figure 3: Expected value of $(1+U)^{\sigma}$ for different numbers of unique values $k=k_{n}$, as the number of observations $n$ increases. The discount parameter is fixed to $\sigma=0.3$.
Figure 4: Left: Unique values $k=k_{n}$ for a conditionally i.i.d. sample from a Dirichlet process of parameters $(\bar{\alpha}=1,\bar{P}_{0}=\mathcal{N}(0,1))$, as the number of observations $n$ increases. Right: $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))$ when $\tilde{\mu}^{1}$ is a generalized gamma$(\alpha_{1}=100,P_{0}^{1}=\mathcal{N}(0,1),\sigma)$ CRM and $\tilde{\mu}^{2}$ is a gamma$(\alpha_{2}=100,P_{0}^{2}=\mathcal{N}(0,1))$ CRM, as the number of observations $n$ increases exponentially. The observations are conditionally i.i.d. sample from a Dirichlet process of parameters $(\bar{\alpha}=1,\bar{P}_{0}=\mathcal{N}(0,1))$.
Figure 5: Left: Unique values $k=k_{n}$ for a conditionally i.i.d. sample from a Pitman-Yor process of parameters $(\bar{\alpha}=1,\bar{\sigma}=0.9,\bar{P}_{0}=\mathcal{N}(0,1))$, as the number of observations $n$ increases. Right: $\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))$ when $\tilde{\mu}^{1}$ is a generalized gamma$(\alpha_{1}=100,P_{0}^{1}=\mathcal{N}(0,1),\sigma)$ CRM and $\tilde{\mu}^{2}$ is a gamma$(\alpha_{2}=100,P_{0}^{2}=\mathcal{N}(0,1))$ CRM, as the number of observations $n$ increases exponentially. The observations are conditionally i.i.d. sample from a Pitman-Yor process of parameters $(\bar{\alpha}=1,\bar{\sigma}=0.9,\bar{P}_{0}=\mathcal{N}(0,1))$.
Theorem 19.
Let $\tilde{\mu}_{1},\tilde{\mu}_{2}$ as in Proposition 17 and let $(x_{n})_{n\geq 1}$ be s.t. $k_{n}\ll n$. Then as $n\rightarrow+\infty$,
$\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))\asymp\max(n^{-1/(1+\sigma)},k/n).$
In particular, $\mathcal{W}_{\mathrm{BL}}(\mathcal{L}(\tilde{\mu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\mu}^{2*}_{\mathcal{S}}))\lesssim\max(n^{-1/(1+\sigma)},k/n)$.
Theorem 19 shows many interesting properties of the merging between a gamma and a generalized gamma CRM. First, recall that when comparing two gamma CRMs as in Section 5.1, the change in the parameters $(\alpha,P_{0})$ did not change the merging rate of $1/n$. In contrast, Theorem 19 shows that when the number of distinct values $k$ diverges sufficiently slowly, a change in $\sigma$ has a heavy impact on the merging, whose rate of convergence slows down to $n^{-\frac{1}{1+\sigma}}\in(1/n,1/\sqrt{n})$. To illustrate this situation, in Figure 4 we sampled conditionally i.i.d. observations from a Dirichlet process of parameters $(\bar{\alpha},\bar{P}_{0})$, since it is well known that $k_{n}\sim\bar{\alpha}\log(n)$ a.s. (Korwar and Hollander, 1973) and thus $k_{n}\lesssim n^{\frac{\sigma}{1+\sigma}}$ a.s. for every $\sigma$.
Moreover, Theorem 19 shows that as the number of distinct values $k$ increases, the rate drastically deteriorates and does not depend anymore on $\sigma$. This is pictured in Figure 5, where we sampled conditionally i.i.d. observations from a Pitman-Yor process (Pitman and Yor, 1997) of parameters $(\bar{\alpha},\bar{\sigma},\bar{P}_{0})$, since in such a case it is well known that $\mathbb{E}(k_{n})\asymp n^{\bar{\sigma}}$ (Pitman, 2006), and thus for $\bar{\sigma}\in[0.5,1)$ we expect $k_{n}\gtrsim n^{\frac{\sigma}{1+\sigma}}$ for every $\sigma$.
Eventually, if $k\asymp n$ there will be no merging of opinions. As a notable example, this happens a.s. when the observations are sampled i.i.d. from a continuous distribution, as described in the next result. It is worth underlining that the lack of merging should not worry the practitioner too much: since model (13) with normalization gives a positive probability to ties among the observations, i.e., $\mathbb{P}(X_{i}=X_{j})>0$, when dealing with independent and continuous data other models for random densities are to be preferred, such as kernel mixtures over a random probability measure.
Proposition 20.
Let $\tilde{\mu}_{1},\tilde{\mu}_{2}$ as in Proposition 17 and let $x_{n}\stackrel{{\scriptstyle\mbox{\scriptsize{\rm iid}}}}{{\sim}}P$, where $P$ is a continuous distribution. Then as $n\rightarrow+\infty$, for a.e. realization of $(x_{n})_{n\geq 1}$,
$\mathcal{W}_{\mathrm{AD}}(\mathcal{L}(\tilde{\nu}^{1*}_{\mathcal{S}}),\mathcal{L}(\tilde{\nu}^{2*}_{\mathcal{S}}))\rightarrow\sigma\bigg{(}\int_{0}^{+\infty}\bigg{|}\frac{\Gamma(-\sigma,t)}{\Gamma(1-\sigma)}-\Gamma(0,t)\bigg{|}\mathrm{d}t+\mathcal{W}_{d_{\mathbb{X},\text{t}}}(P_{0},P)\bigg{)}.$
Moreover, the limit is increasing in $\sigma$.
We observe that the limit in Proposition 20 does not depend on $\alpha$ and it is composed by two terms. The first term coincides with the distance a priori between a scaled gamma and scaled generalized gamma with same total base measure, and can thus be seen as a measure of discrepancy between the priors for the jumps. The second term coincides with the distance between the base measure and the law of the observations, and can thus be seen as a measure of discrepancy between the (common) prior for the atoms and the data. As expected, the limit vanishes as $\sigma\to 0$ and it increases with $\sigma$.
From a statistical perspective, Theorem 19 detects the regimes where using a normalized generalized gamma CRM will actually guarantee a different learning outcome, with the following take-home message: either the number of observations $n$ is sufficiently small or the number of distinct values $k$ should be sufficiently large, and how large it must be depends on the choice of discount $\sigma$, namely, $k>n^{\frac{\sigma}{1+\sigma}}$. As notable examples, this happens when the observations are i.i.d. from a continuous distribution or when they are conditionally i.i.d. given a normalized generalized gamma CRM or a Pitman-Yor process, but it does not happen when the observations are i.i.d. from a discrete distribution with a finite number of atoms or when they are conditionally i.i.d. given a Dirichlet process or a hierarchical Dirichlet process.
Acknowledgments
The authors would like to thank A. Lijoi and I. Prünster for valuable insights. Part of this research has been done during a visit of the second author to the Department of Statistics at the University of Warwick, which is warmly acknowledged for the hospitality. The second author is also affiliated to the Bocconi Institute for Data Science and Analytics (BIDSA). M. Catalano was partially supported by a Research Development Fund from the University of Warwick.
Supplementary material
Contains the proofs of the results of the present work.